OpenStudy (anonymous):

help please im on my last question and i keep doing something wrong ): –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

start with \[-2x+6\geq 5-x-3\]

OpenStudy (anonymous):

then combine like terms on the right to get \[-2x+6\geq -x+2\] add \(2x\) to both sides and get \[6\geq x+2\] and finish by subtracting 2 from both sides

OpenStudy (anonymous):

okay the right side was what was confusing me i get how to solve it now thankyou!(: @satellite73